OpenStudy (anonymous):

Evaluate w/o using a calulator! tan (-3 pi/4) Show work or explain it please! Thanks

OpenStudy (anonymous):

1. tan(-x) = -tan(x) (from drawing)  therefore tan(-3pi/4) = -tan(3pi/4) 2. convert to degrees to make things simpler 3pi/4 radians =135 degrees 3. tan(180 - x) = -tan(x) so -tan(135) = -(-tan(45)) 4. tan(45) =1 so -(-tan(45)) = -(-1) = 1